OpenStudy (anonymous):

how can you resolve this pro: 3(4-x)=5-x+ 6(2-x)

OpenStudy (anonymous):

All you need to do is distribute the multiplication, then group together like terms. $$ 3(4-x) \Rightarrow 12-3x \\ 6(2-x) \Rightarrow 12 - 6x \\ \Rightarrow 12-3x = 5-x+12-6x\\ \Rightarrow 12-5-12 = 3x-x-6x\\ \Rightarrow -5 = -4x\\ \Rightarrow x = \frac{5}{4} $$

OpenStudy (anonymous):

\[3(4-x) \Rightarrow 12-3x\\ 6(2-x) \Rightarrow 12-6x\\ \Rightarrow 12-3x = 5-x+12-6x\\ \Rightarrow 12-5-12 = 3x-x-6x\\ \Rightarrow -5 = -4x \\ \Rightarrow x = \frac{5}{4}\]

OpenStudy (shadowfiend):

Let's try this: \[\begin{align} &3(4-x) \Rightarrow 12-3x\\ &6(2-x) \Rightarrow 12-6x\\ &\Rightarrow 12-3x = 5-x+12-6x\\ &\Rightarrow 12-5-12 = 3x-x-6x\\ &\Rightarrow -5 = -4x \\ &\Rightarrow x = \frac{5}{4} \end{align}\]